# WiLLM: An Open Wireless LLM Communication System

### v1

@@ Abstract.
1. Introduction
The convergence of artificial intelligence and wireless communications represents one of the most significant technological intersections of the current decade. As computational capabilities and algorithmic sophistication continue to advance, LLMs have emerged as transformative tools with far-reaching implications across diverse application domains. The integration of these models into wireless communication systems has catalyzed a paradigm shift, enabling advanced intelligent services ranging from real-time multilingual translation to context-aware augmented reality applications (Bariah et al., 2024). As this integration accelerates, telecommunication infrastructure faces unprecedented challenges in supporting the computational demands of these resource-intensive models. Current deployments predominantly rely on either edge computing paradigms or centralized cloud processing approaches, each presenting significant limitations. Edge deployment minimizes transmission latency but suffers from constrained computational capacity and thermal limitations in user equipment (UE), while cloud-based solutions introduce unpredictable internet latency that severely impacts quality-of-experience (QoE) for interactive LLM services (Zheng et al., 2024). Moreover, the exponential growth in concurrent LLM service requests threatens to overwhelm both communication channels and computational resources, creating critical bottlenecks in service provisioning.
Recent implementations across several institutions in China, including deployments of Deepseek at multiple universities (Engineer, 2025), demonstrate the feasibility of local LLM deployments at the network edge (Boateng et al., 2024). While these edge deployments represent a promising step toward LLM integration within telecommunications infrastructure, they still encounter resource limitations and scalability constraints inherent to edge computing. These implementations nonetheless validate the technical viability of localized LLM services and establish a foundation for our proposed migration of LLM computational resources to the core network, where greater processing capacity and more efficient resource allocation can be leveraged. This transition from edge to core network deployment represents a natural evolution in the LLM service architecture, addressing the scalability limitations of edge deployments while maintaining the latency advantages over cloud-based solutions (Chaoub and Elkotob, 2025).
The current network architecture exhibits three fundamental drawbacks hindering effective LLM service integration: (1) the absence of dedicated open testbeds for LLM-specific communication scenarios, (2) the inherent conflict between centralized computation paradigms and distributed network architectures, and (3) the lack of viable business models for telecom operators in LLM service provisioning. Existing network slicing architectures provide potential foundations for addressing these challenges through resource isolation and service customization. However, conventional slicing methodologies primarily focus on traditional communication services and lack the specialized mechanisms required for computationally intensive LLM workloads (Dehkordi et al., 2023). Furthermore, the absence of standardized APIs and resource management frameworks inhibits the development of cross-layer optimization strategies essential for efficient LLM service delivery.
Another critical limitation is the high usability barrier of current systems. Existing platforms require specialized wireless expertise and native slicing support on user equipment, restricting access primarily to communication specialists (Zheng et al., 2024). This complexity explains why testbeds are often used only by their developer communities, significantly limiting cross-disciplinary research and broader adoption for LLM service experimentation. Moreover, the field lacks comprehensive empirical datasets capturing cross-layer performance metrics, standardized benchmarking methodologies for systematic comparison, and representative case studies demonstrating viability in resource-constrained environments. This empirical deficit perpetuates the disconnect between algorithmic innovation and operational deployment, hampering the development of generalizable principles for wireless LLM communication systems.
Motivated by the above system and networking challenges, we present WiLLM, the world’s first open-source wireless communication system specifically designed for mobile LLM services. Building upon the solid foundation provided by OpenAirInterface’s modular architecture, WiLLM strategically extends and enhances this established framework with specialized components for LLM service integration, creating a system that maintains compatibility with existing infrastructure while introducing transformative capabilities for AI-driven communications research.
The work unifies the following progressive contributions:
1.1. Paradigm
Novel Network Paradigm for LLM Services (Section 3). WiLLM establishes a novel paradigm by deploying LLMs in core networks (CNs) with GPU infrastructure, enabling distributed LLM inference services strategically positioned at the convergence point between high-bandwidth backbone connections and cellular network edges. This architectural configuration optimizes computational-communication tradeoffs while facilitating operator monetization through subscription-based network slice allocation.
1.2. Architecture
“Tree-Branch-Fruit” Slicing Architecture (Section 3.3). We propose a multi-layered slicing architecture wherein the “tree” represents foundational network infrastructure, the “branches” correspond to conventional communication service slices, and the “fruits” denote specialized LLM service slices. This architecture establishes a “slice-as-LLM-service” paradigm while maintaining compatibility with existing slicing mechanisms, achieving coordinated scheduling across multiple users and slices through cross-layer resource optimization.
1.3. System
The Wireless Communication System Specialized for LLM Service (Section 4). Based on OpenAirInterface, WiLLM implements five pivotal innovations: dynamic slice compatibility, universal UE compatibility, dual-mode resource scheduling, cross-layer API framework, and flexible deployment support. This integrated solution addresses critical limitations in existing wireless infrastructure while preserving backward compatibility.

### v2

@@ 1. Introduction
Figure 2. The LLM Services Pipeline over WiLLM (Liu et al., 2024).
The challenge extends beyond simple capacity constraints. Our measurements across diverse deployment scenarios reveal that LLM services violate core assumptions underlying current wireless system designs. Traditional traffic engineering approaches assume predictable patterns, unidirectional dominance, and static computational requirements (Andrade and Wickboldt, 2025). LLM services break these assumptions through their generative nature, creating workloads that dynamically shift between computation and communication bottlenecks within single sessions. This fundamental mismatch prevents effective service delivery and optimization using existing tools and methodologies.
Compounding these technical challenges, the research ecosystem lacks accessible platforms for studying LLM behavior in wireless environments. Current open source testbeds such as OpenAirInterface (Nikaein et al., 2014) and srsRAN (Gomez-Miguelez et al., 2016) require deep telecommunications expertise and specialized hardware, effectively excluding general researchers and developers from wireless experimentation. Meanwhile, LLM deployment frameworks like Ollama (Ollama, 2025) and TensorRT-LLM (NVIDIA, 2025) operate without awareness of wireless constraints. This divide between communities impedes innovation at their intersection, where the most impactful advances will emerge in mobile AI services (Zheng et al., 2024).
Our extensive empirical analysis, detailed in Section 2, identifies specific characteristics that distinguish LLM services from traditional applications and motivate specialized system design. These findings drive WiLLM, an open-source platform that bridges the gap between wireless infrastructure and LLM service. Rather than treating LLM services as generic data traffic, WiLLM incorporates architectural innovations specifically addressing their unique requirements while maintaining accessibility for researchers without wireless expertise.
Figure 1 and 2 illustrates WiLLM’s advantages and the service pipeline. Built on OpenAirInterface’s robust protocol implementation, WiLLM extends the base platform with LLM-aware capabilities while preserving compatibility with standard devices and networks. The system enables researchers to experiment with real cellular networks using commercial hardware, deploy various LLM models from Llama (ggml-org, 2025) to LLaVA (Liu et al., 2023), and observe detailed cross-layer interactions through comprehensive monitoring interfaces. These capabilities transform wireless LLM research from a specialized domain requiring extensive telecommunications knowledge to an accessible field where AI researchers can directly contribute innovations. To evaluate the platform’s capability, we conducted a case study with smart glasses, where a team with minimal wireless expertise successfully deployed and optimized a multimodal LLM service. This validation confirms that WiLLM achieves its dual objectives: providing the technical capabilities necessary for LLM service delivery while removing barriers that have historically separated wireless and AI research communities.
This paper makes three primary contributions. (1) We provide empirical evidence establishing how LLM services differ fundamentally from traditional network applications, based on analysis of over 1.6 million synchronized measurements across network layers. (2) We present WiLLM as an open experimental platform that democratizes wireless LLM research through architectural innovations and intuitive interfaces. (3) We release our complete implementation, comprehensive dataset, and benchmarking tools to foster an ecosystem for wireless LLM innovation.
The remainder of this paper is organized as follows. Section 2 analyzes the unique characteristics of LLM services through empirical measurements. Section 3 presents WiLLM’s system architecture and design principles. Section 4 details our implementation on OpenAirInterface. Section 5 validates our design through comprehensive evaluation. Section 6 reviews related work, and Section 7 concludes.
2. LLM Services: Why Different from Traditional DNN over Wireless Systems?
We deployed WiLLM to systematically characterize LLM service behavior across commercial UE devices, a 5G RAN with configurable resource scheduling, and a GPU-accelerated Core Network. Our experimental design combines static and dynamic configurations for both UE and gNB components across four operational scenarios, isolating the effects of system adaptability and network slicing on service performance. Multimodal workloads include image-to-text queries that stress uplink capacity, and text-to-image generation that creates diverse downlink requirements. NTP-based synchronization achieves millisecond-level accuracy for cross-layer correlation, enabling us to trace requests from application initiation through network transmission to inference completion. Through careful control of network conditions and computational isolation, we collected extensive empirical evidence revealing how LLM services fundamentally differ from traditional wireless applications, as detailed in the following three insights.
2.1. Insight 1: Bidirectional Heavy Load Asymmetry
Traditional DNN applications in wireless networks exhibit clear unidirectional patterns. For example, image classification services upload raw images and receive lightweight labels, while video analytics applications stream uplink data for cloud processing. This unidirectional dominance enables straightforward optimization: allocate resources primarily for the dominant direction.
LLM services fundamentally break this pattern due to their multimodal interaction capabilities. Within a single session, users seamlessly transition between uploading images for scene understanding and requesting image generation from text descriptions. Our measurements reveal that this creates two contrasting operational modes with opposite bottlenecks. As shown in Figure 3, in uplink-heavy scenarios (image-to-text), computational inference at the LLM dominates the end-to-end latency, accounting for 74% to 87% of total delay, while network transmission contributes marginally. Conversely, Figure 4 demonstrates that in downlink-heavy scenarios (text-to-image), network transmission becomes the primary constraint, consuming 81% to 86% of total latency, with inference time reduced from 12% to 17%.
This asymmetry stems from the fundamental nature of multimodal LLMs. Processing high-resolution images requires extensive computation for feature extraction and semantic understanding, creating uplink inference bottlenecks. Meanwhile, generating detailed visual content produces large data payloads that stress downlink capacity. The continuous nature of LLM interactions means both patterns occur within the same session, making traditional unidirectional optimization strategies ineffective. Systems must dynamically recognize and adapt to these contrasting requirements, a capability absent in current wireless architectures designed for predictable traffic patterns.
2.2. Insight 2: Computation-Communication Coupling in Slicing
Traditional DNN workloads demonstrate predictable computational patterns. A ResNet-50 (Kaiming He and Sun, 2015) model requires approximately 4 GFLOPs per image regardless of network conditions. A YOLO (Redmon et al., 2016) detector maintains consistent inference time across deployments. This computational predictability decouples inference from network resource allocation, enabling independent optimization of computation and communication resources.
Figure 3. Component distribution and latency analysis for uplink transmissions across resolution groups. The x-axis labels R1-R6 represent various resolution groups for image requests sent to LLM, where R1-R6 correspond to increasing resolutions. The left panel shows the latency composition of different components (uplink, downlink, inference), the center panel presents statistical trends with standard errors, and the right panel quantifies each component’s percentage contribution to total latency.
Figure 4. Component distribution and latency analysis for downlink transmissions across resolution groups. Similar to Figure 3, it provides three complementary views: component distribution, statistical latency metrics, and proportional component analysis for downlink scenarios.
Figure 5. Component distribution and latency analysis across slice groups. Communication resources for uplink slices in Slice 1 to Slice 3 are gradually increasing. The middle panel statistically validates performance differences, while the rightmost panel quantifies how slice configurations shift processing time proportions among uplink, downlink, and inference operations.
LLM services exhibit fundamentally different behavior where computational demands vary dramatically and interact complexly with network resources. Unlike traditional DNNs with fixed architectures processing fixed-size inputs, LLMs generate variable-length outputs through autoregressive decoding. The inference time depends on multiple dynamic factors, such as the input and output token count, sampling temperature, beam search width, accumulated context length, and more. Our measurements reveal that identical prompts can result in inference times varying by over 300% based on these generation parameters.
More critically, this computational variability interacts with network slicing to create dynamic bottleneck migration. As shown in Figure 5, when network slice resources increase from minimal to abundant allocation, the system bottleneck shifts dramatically. Under constrained slice configurations, inference time dominates at 59.6% of total latency, as limited network resources throttle the overall pipeline. However, with enhanced slice resource allocation, the bottleneck migrates to transmission, which grows to 54.7% of latency. This migration occurs because improved network resources enable faster data transfer, exposing computational limitations that were previously masked by network constraints.
This coupling between computation and communication invalidates traditional slice optimization approaches that assume fixed service characteristics. LLM services require dynamic adaptation mechanisms that consider both computational state and network conditions, adjusting resource allocation based on real-time bottleneck identification rather than static service profiles.
2.3. Insight 3: Token Stream Unpredictability
Traditional DNN applications generate highly predictable network traffic. Classification models produce fixed-size outputs (Krizhevsky et al., 2017); object detectors generate bounded detection lists (Zhao et al., 2019; Redmon et al., 2016); even video streams maintain consistent bitrates through rate control (Om et al., 2022; Alsmirat et al., 2023). This predictability enables efficient resource allocation through statistical multiplexing and quality-of-service guarantees based on well-characterized traffic models.
LLM token streams violate these assumptions through three unique characteristics (Stojkovic et al., 2025). First, the generative nature creates inherent unpredictability in output length. A simple query may yield trigger a terse response or an elaborate explanation, depending on subtle prompt variations or sampling randomness. Second, token generation exhibits extreme burstiness as the model alternates between rapid token transmission and computational pauses for attention calculation. Third, multi-turn interactions introduce state dependencies where previous exchanges influence subsequent resource requirements through context accumulation.
Figure 6. Temporal analysis of PRB allocation patterns across three implementation scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom). The x-axis represents timestamps in microseconds, while the y-axis denotes current resource blocks allocated to the UE. The visualization demonstrates the effectiveness of the slicing architecture in providing differentiated resource allocation.
Figure 7. Temporal analysis of actual byte transmission across three scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom). Despite controlled PRB allocation, actual byte transmission shows high variance, particularly pronounced in LLM traffic patterns.
Our measurements, visualized in Figures 6 and 7, reveal the practical implications of this unpredictability. Despite controlled Physical Resource Block (PRB) allocation following defined slice policies, actual byte transmission shows high variance and non-linear relationships with allocated resources. This non-linearity is particularly pronounced in LLM traffic compared to conventional data flows, as the stochastic generation process creates transmission bursts that poorly align with periodic resource scheduling.
The unpredictability extends beyond simple variance. Token streams exhibit complex temporal correlations as the model generates semantically coherent sequences, creating traffic patterns that defy traditional Markovian models. Resource allocation strategies based on average behavior or peak provisioning both prove inadequate: average-based allocation suffers from frequent deadline violations during generation bursts, while peak provisioning wastes resources during computation-bound phases. Effective LLM service delivery requires fine-grained monitoring and adaptive resource management that responds to real-time generation patterns rather than statistical traffic models.
These insights reveal why LLM services demand new wireless designs: bidirectional asymmetry requires direction-aware scheduling; computation-communication coupling needs dynamic slice optimization; and token stream unpredictability calls for adaptive resource management. Unlike traditional DNN applications, these characteristics necessitate rethinking wireless architectures. The following section presents WiLLM, our platform addressing these challenges through targeted design principles and implementation choices.
3. WiLLM System Architecture
The insights from Section 2 drive WiLLM’s architectural decisions. This section presents our design philosophy, system overview, and key innovations that collectively enable efficient LLM service delivery over wireless networks while democratizing access for researchers without wireless expertise.
3.1. Design Philosophy
WiLLM’s design philosophy centers on making LLM services accessible to researchers while addressing their unique technical challenges. Based on our empirical insights and the limitations of existing systems, we establish three core principles that guide our architectural decisions.
Principle 1: Universal Accessibility. Unlike existing testbeds that require specialized devices with native slicing support (Saboorian and Xiang, 2017), WiLLM must enable any standard device to leverage network slicing capabilities for fine-grained resource allocation. This principle drives our universal UE compatibility design through application-layer tunneling, allowing commodity devices to access slice-specific resources that guarantee LLM service requirements. By bridging the gap between slice-unaware devices and slice-enabled networks, we empower standard smartphones and laptops to benefit from differentiated QoS provisioning, ensuring predictable latency and bandwidth allocation crucial for interactive LLM applications. This democratizes advanced network resource management beyond specialized research equipment.
Principle 2: Dynamic Adaptability. The complex interplay between LLM computational variability and network conditions demands systems that can adapt in real-time. Static configurations fail to address the migrating bottlenecks and changing resource requirements inherent in LLM services. This principle motivates our dynamic slice compatibility mechanisms, multi-UE multi-slice scheduling capabilities, and dual-mode resource allocation, enabling the system to respond to varying workload characteristics without manual reconfiguration.
Principle 3: Cross-Layer Observability. The tight coupling between application-layer token generation, network-layer transmission, and computation-layer inference requires visibility and control across all system layers. Traditional layered architectures that hide internal state prevent the coordinated optimization necessary for LLM services. Real-time adaptation to LLM’s dynamic behavior requires understanding not just what happens at each layer, but how events cascade across layers to impact end-to-end performance. This principle drives our comprehensive cross-layer API framework and monitoring infrastructure, exposing system state while maintaining appropriate abstractions for non-experts.
These principles balance technical requirements from our measurements with practical research needs. Rather than prescribing specific algorithms, WiLLM provides an open platform for deploying and evaluating arbitrary scheduling strategies, removing barriers that have limited innovation at the intersection of wireless networks and LLM services.
3.2. Key System Innovations
3.2.1. Core Network GPU Deployment
WiLLM fundamentally reimagines LLM service deployment by positioning GPU resources within the core network rather than at edge nodes or remote clouds. As illustrated in Figure 8, this architectural choice addresses the unique challenges of LLM services while leveraging existing telecommunications infrastructure.
Traditional cloud deployments (left side of Figure 8) require traffic to traverse the public internet, introducing variable latency, unpredictable congestion, and loss of operator control over service quality (Chen et al., 2024). Edge deployments attempt to reduce latency but face severe limitations: insufficient computational capacity for large language models, thermal constraints in compact edge nodes, and coordination challenges across distributed edge sites (He et al., 2024; Cai et al., 2024; Nelson and Toloui, 2023; Tang et al., 2023).
Our core network deployment (right side of Figure 8) positions GPU clusters at the convergence point between high-capacity backhaul networks and the radio access network. This location offers unique advantages for LLM services. First, it provides access to substantial computational resources through dedicated GPU infrastructure connected via high-speed interconnects. Second, it maintains low latency by avoiding internet traversal while serving multiple base stations from a centralized location. Third, it enables coordinated resource management as both communication and computation resources fall under unified administrative control.
This deployment paradigm particularly benefits from the bidirectional asymmetry of LLM services. During uplink-heavy inference, the proximity to radio networks minimizes data transfer delays while powerful GPUs handle computational bottlenecks. For downlink-heavy generation, the high-bandwidth backhaul connections efficiently distribute generated content. The centralized nature also facilitates resource sharing across multiple concurrent LLM sessions, improving utilization compared to distributed edge deployments.
3.2.2. Tree-Branch-Fruit Slicing Architecture
Figure 9. Tree-Branch-Fruit network slicing architecture with hierarchical service organization. The tree represents base infrastructure, branches correspond to traditional service categories, and fruits denote specialized LLM services with differentiated pricing models.
Building upon core network GPU deployment, WiLLM introduces the Tree-Branch-Fruit slicing architecture that extends traditional network slicing for LLM service differentiation. As depicted in Figure 9, this three-tier hierarchy enables flexible service provisioning while maintaining compatibility with existing 5G slicing mechanisms.
The Tree layer represents the foundational radio access network infrastructure shared across all services. This includes the physical base stations (gNBs), spectrum resources, and basic connectivity functions. The tree provides common entry points for all traffic while implementing isolation mechanisms to prevent interference between different service branches.
The Branch layer corresponds to traditional 5G network slices: enhanced Mobile Broadband (eMBB), Ultra-Reliable Low-Latency Communication (URLLC), and massive Machine Type Communication (mMTC). These branches maintain standard 3GPP-defined behaviors, ensuring backward compatibility with existing devices and services. Each branch implements specific quality-of-service parameters, resource allocation policies, and isolation guarantees as specified in 5G standards.
The Fruit layer introduces LLM-specific service slices that extend from appropriate branches based on their primary characteristics. Each fruit slice is configured for particular LLM applications with tailored resource allocations, model selections, and pricing tiers. For instance, a real-time translation service might extend from the URLLC branch with a lightweight model ensuring low latency, while a creative writing assistant could branch from eMBB with a larger model prioritizing quality over speed.
This hierarchical architecture enables several key capabilities. Service differentiation allows operators to offer multiple LLM service tiers with varying performance guarantees and pricing models, as indicated by the monetary symbols in Figure 9. Dynamic fruit creation permits rapid deployment of new LLM services without modifying core network functions. Resource isolation ensures that computational or communication demands from one LLM service do not impact others. Most importantly, the architecture maintains a clear evolutionary path where existing networks can progressively add LLM capabilities through fruit slice deployment.
Figure 10. Comprehensive system architecture of WiLLM with UE, gNB, CN, and LLM Server Components. The diagram illustrates data flows, control pathways, and the integration of LLM-specific components within the telecommunications infrastructure.
WiLLM implements five key innovations that collectively address the unique challenges of LLM service delivery while maintaining accessibility for researchers, as illustrated in the comprehensive system architecture shown in Figure 10.
3.2.3. Dynamic Slice Compatibility
Traditional network slicing assumes static service characteristics (Andrade and Wickboldt, 2025), but LLM services exhibit highly variable resource requirements based on model selection, generation parameters, and user interactions. WiLLM’s dynamic slice compatibility extends conventional slicing with runtime adaptation capabilities.
As shown in Figure 10, the Slice Manager components in both gNB and UE subsystems implement bidirectional communication through UE Update and UE State Report pathways. This enables real-time slice reconfiguration based on observed performance metrics. The Branch Slice Policy Executor and Fruit Slice Policy Executor translate high-level service requirements into concrete resource allocation decisions, adapting to changing computational loads and network conditions.
The dynamic compatibility specifically addresses LLM’s variable inference times and unpredictable token generation patterns. When the system detects a shift from computation-bound to communication-bound operation, slice parameters automatically adjust to maintain service quality. This adaptation occurs transparently, requiring no manual intervention from researchers.
3.2.4. Universal UE Compatibility
A critical barrier in existing platforms is the requirement for specialized UE equipment with native slicing support. WiLLM eliminates this barrier through application-layer tunneling mechanisms that enable any standard device to access LLM services.
The Slice Manager of UE component in Figure 10 implements a middleware layer that intercepts application traffic and encapsulates it with appropriate slice identifiers. This approach bypasses the limitations of current UE implementations, which often lack complete slicing support. The Configuration Manager and Hot Start Module work together to establish slice contexts without requiring protocol-level modifications.
This innovation is crucial for democratizing access to LLM service research. Researchers can use commercial smartphones, laptops, or IoT devices without specialized firmware or hardware modifications. The application-layer approach also enables rapid prototyping of new service types without waiting for standardization or device updates.
3.2.5. Multi-UE Multi-Slice Scheduling
LLM services create complex scheduling challenges when multiple users with different service requirements share network resources. WiLLM implements sophisticated scheduling mechanisms that coordinate resources across multiple UEs and slice types simultaneously.
The Scheduler module at the center of Figure 10 receives inputs from multiple sources: Buffer Status from the Buffer Manager, Retransmission Requirements from the HARQ Manager, and Slice Allocation policies from the mapping components. This multi-dimensional optimization problem is solved through a hierarchical approach: first allocating resources across slices based on their aggregate requirements, then scheduling individual UEs within each slice based on service-specific metrics.
The scheduling algorithm specifically accounts for LLM characteristics. Token generation bursts receive priority handling to minimize user-perceived latency, while background model updates are scheduled during resource valleys. The system maintains fairness across users while respecting the differentiated service levels promised by different fruit slices.
3.2.6. Dual-Mode Resource Scheduling
To balance implementation efficiency with optimization sophistication, WiLLM implements complementary embedded and separated scheduling modes, each suited to different operational scenarios.
In embedded mode, slice-aware resource allocation integrates directly within the main scheduling pipeline. This provides low-latency adaptation suitable for real-time LLM interactions. The scheduler modifies its decisions based on slice policies without external intervention, enabling microsecond-scale adaptations to changing conditions.
In separated mode, an external optimization engine analyzes longer-term patterns and computes globally optimal resource allocations. This engine interfaces through the Resource Update pathway, enabling sophisticated algorithms including machine learning-based prediction and multi-objective optimization. The separated mode excels at identifying persistent patterns in LLM usage and pre-allocating resources accordingly.
The dual-mode design allows researchers to experiment with different scheduling strategies without modifying core system components. Simple policies can be implemented in embedded mode for immediate testing, while complex algorithms can be developed in separated mode with full access to historical data and predictive models.
3.2.7. Cross-Layer API Framework
To enable effective research without requiring wireless expertise, WiLLM provides comprehensive APIs that abstract network complexity while exposing necessary control and monitoring capabilities. As shown in the online Figure 16, the API architecture implements a three-tier structure.
The User Management API handles service registration, authentication, and preference management. Researchers interact with familiar application-level concepts like ”model selection” and ”latency requirements” rather than radio network parameters. The API automatically translates these high-level specifications into appropriate network configurations.
The System Management API provides slice lifecycle management and service orchestration. Through RESTful interfaces, applications can discover available services, request specific configurations, and monitor performance metrics. The API includes WebSocket endpoints for real-time monitoring of token generation rates, latency breakdowns, and resource utilization.
The Resource Management API offers fine-grained control for advanced users who need to optimize specific aspects of service delivery. This includes interfaces for custom scheduling policies, resource reservation, and predictive allocation based on application hints. Importantly, this API remains optional—basic LLM services function without any low-level configuration.
The cross-layer design ensures information flows efficiently between system components. Token generation patterns from the application layer inform network scheduling decisions, while network congestion indicators trigger application-layer adaptation. This bidirectional information flow, illustrated by the various feedback pathways in Figure 10, enables holistic optimization impossible in traditional layered architectures.
These five innovations work synergistically to create a system that is both powerful enough to handle LLM’s unique challenges and accessible enough for researchers without wireless expertise. The following section details how we realized these concepts through concrete implementation on the OpenAirInterface platform.
4. Implementation
Transforming WiLLM’s architectural vision into a functional system required substantial engineering effort. This section details our implementation approach, focusing on how we extended OpenAirInterface to support LLM services while maintaining compatibility and usability.
4.1. Building on OpenAirInterface
We selected OpenAirInterface (OAI) as our implementation foundation for several compelling reasons. First, OAI provides a complete, modular implementation of 4G/5G protocol stacks from physical layer to core network, enabling end-to-end system modifications. Second, its open-source nature allows deep modifications necessary for LLM-specific optimizations. Third, OAI’s active community and extensive documentation lower the barrier for researchers to understand and extend our work. However, implementing WiLLM required substantial extensions beyond OAI’s baseline capabilities. Our modifications span three critical areas:
Slice Management Extensions. While OAI includes basic network slicing support, it lacks the dynamic adaptation required for LLM services. We extended the Radio Resource Control (RRC) and Service Data Adaptation Protocol (SDAP) layers to support runtime slice reconfiguration. The modifications enable slice parameters to be adjusted based on real-time performance metrics without disrupting ongoing sessions. We implemented new signaling procedures that propagate slice updates from the core network through the RAN to individual UEs, maintaining consistency across the system.
Scheduler Enhancements. The OAI scheduler originally implements proportional fair and round-robin algorithms optimized for traditional traffic patterns. We augmented it with LLM-aware scheduling logic that recognizes token generation patterns and adapts resource allocation accordingly. Our modifications include: (1) traffic classification modules that identify LLM flows based on packet patterns, (2) dual-mode scheduling support with hooks for external optimization engines, and (3) direction-aware resource allocation that treats uplink and downlink asymmetrically based on identified bottlenecks.
Cross-Layer Information Flow. Traditional OAI maintains strict layer separation, preventing the coordinated optimization necessary for LLM services. We implemented information conduits that enable controlled cross-layer communication while preserving architectural modularity. New interfaces expose internal state from MAC, RLC, and PDCP layers to higher-level orchestration components. Similarly, application-layer hints about token generation patterns flow down to influence scheduling decisions. These modifications required careful design to avoid violating layer abstractions while enabling necessary information exchange.
4.2. Engineering Challenges and Solutions
Implementing WiLLM’s architectural innovations presented several significant engineering challenges that required novel solutions based on the unique characteristics of LLM services identified in Section 2.
Challenge 1: Universal UE Compatibility Without Protocol Modifications. A critical barrier in existing platforms is the requirement for specialized UE equipment with native slicing support. Current community implementations of network slicing remain incomplete and inconsistently supported across device ecosystems, effectively excluding standard commercial devices from accessing advanced network features.
Why existing systems fail: Current platforms like OAI and srsRAN hard-code slice support at the protocol level, requiring modifications to UE protocol stacks for slice access. This approach excludes the vast majority of commercial devices and creates an impractical barrier for researchers who lack the capability to modify device firmware or develop custom UE implementations.
Solution: We developed a middleware-based slice access mechanism through application-layer tunneling techniques. The Slice Manager of UE component implements a compatibility layer that encapsulates LLM service traffic within standard application protocols, enabling infrastructure compatibility while maintaining architectural autonomy. This approach provides telecommunications operators with implementation flexibility independent of underlying radio network slice capabilities, allowing fruit slices to evolve asynchronously from hardware upgrade cycles or standardization timelines.
Challenge 2: Dynamic Slice Reconfiguration for Variable LLM Workloads. LLM services exhibit highly variable resource requirements based on model selection, generation parameters, and user interactions, as demonstrated in our empirical analysis. Traditional network slicing assumes static service characteristics, creating a fundamental mismatch with LLM’s dynamic nature.
Why existing systems fail: Traditional slicing implementations in platforms like FlexRIC and Open5GS follow 3GPP specifications that assume static service profiles. Once configured, slice parameters remain fixed throughout the session lifetime. This rigid approach cannot accommodate LLM’s dynamic resource requirements that vary by orders of magnitude within a single session.
Solution: We extended conventional slicing with runtime adaptation capabilities through bidirectional communication pathways (UE Update and UE State Report) that enable real-time slice reconfiguration based on observed performance metrics. The Branch Slice Policy Executor and Fruit Slice Policy Executor translate high-level service requirements into concrete resource allocation decisions, adapting to changing computational loads and network conditions without disrupting ongoing sessions.
Challenge 3: Cross-Layer Coordination for Computation - Communication Coupling. Our measurements revealed tight coupling between LLM computational variability and network performance, where slice configuration changes cause dynamic bottleneck migration between computation and communication. Traditional layered architectures prevent the coordinated optimization necessary for managing this coupling.
Solution: We implemented a comprehensive cross-layer API framework that facilitates vertical integration between application requirements and network-level resource allocation. The system includes performance feedback loops, service-specific optimization pathways that enable LLM Service ID information to influence network decisions, and resource utilization monitoring that provides visibility into consumption patterns across system layers while maintaining appropriate abstractions for non-experts.
Challenge 4: Multi-UE Multi-Slice Scheduling Complexity. LLM services create complex scheduling challenges when multiple users with different service requirements share network resources. The variable inference times and unpredictable token generation patterns identified in our analysis significantly complicate resource allocation compared to traditional traffic.
Solution: We implemented a hierarchical scheduling mechanism that operates through a two-phase process: a global allocation phase distributes available physical resource blocks across active slices according to their relative priorities and resource guarantees, followed by an intra-slice scheduling phase that allocates resources to individual UEs within each slice based on service-specific metrics. This approach ensures fair and efficient resource utilization while maintaining service isolation.
Challenge 5: Temporal Synchronization Across Distributed Components. Correlating LLM service events across UE, RAN, and CN components requires precise time synchronization, particularly challenging given the variable processing delays inherent in LLM inference and the need for microsecond-level accuracy in performance analysis.
Solution: We implemented NTP-based distributed synchronization across all system components, calibrating devices through a common server to achieve microsecond-level sampling precision. By recording timestamps at both UE and server endpoints and implementing latency compensation algorithms, we maintained synchronization errors within ±1.0 milliseconds. This precision enables the accurate cross-layer correlation analysis that forms the foundation of our empirical insights.
These engineering solutions demonstrate that WiLLM’s architectural innovations can be realized through practical implementations that address the unique challenges of LLM service delivery while maintaining system stability and research accessibility.
5. WiLLM in Action: Deployment, Dataset, and Case Study
This section demonstrates WiLLM’s practical capabilities through comprehensive deployment, data collection, and real-world application development. Unlike conventional wireless testbeds that focus on protocol compliance, WiLLM enables researchers to study and optimize LLM services across realistic wireless environments.
5.1. Real-World Deployment
WiLLM’s operational deployment demonstrates the system’s capability to deliver end-to-end LLM services while providing a robust experimentation platform for researchers.
System Performance and Capabilities: As illustrated in our online appendix Figure 14, our deployed testbed achieves full end-to-end LLM service delivery with commercial devices. The system successfully handles concurrent multi-user sessions with differentiated service levels through the Tree-Branch-Fruit slicing architecture. The deployment maintains stable operation across diverse LLM workloads, from lightweight text generation to compute-intensive multimodal processing, validating the practical viability of core network GPU deployment for production environments.
Research Platform Features: The testbed provides researchers with unprecedented experimental control through its integrated monitoring and adaptation capabilities. Real-time performance visualization enables immediate observation of cross-layer interactions, while programmable APIs allow dynamic reconfiguration without service interruption. Researchers can experiment with various LLM models (LLaVA, Llama 3.2) and network configurations through simple parameter adjustments, without requiring telecommunications expertise. The platform’s stability and ease of use have enabled extended experimental campaigns lasting multiple days without manual intervention.
Accessibility and Adoption: A key achievement is enabling AI researchers to experiment with communication resource allocation without deep protocol stack expertise. Through intuitive Python APIs and web interfaces, researchers can specify scheduling policies, adjust slice parameters, and control bandwidth allocation—making wireless resource management accessible to those focused on LLM optimization. The platform abstracts complex protocol details while exposing meaningful control knobs: AI researchers can explore how different QoS configurations, priority schemes, and resource partitioning strategies impact their models’ performance. Standard devices connect seamlessly through our application-layer approach, allowing experiments on commodity hardware. This design bridges the gap between AI and networking communities, enabling AI researchers to actively engage with communication-layer optimizations rather than treating the network as a black box.